OpenStudy (anonymous):

The product of two consecutive positive even integers is 12,320. What are the two integers?

Parth (parthkohli):

So, we have x and x + 2 as the two integers. \( \color{Black}{\Rightarrow x(x + 2) = 12320}\) \( \color{Black}{\Rightarrow x^2 + 2x = 12320}\) \( \color{Black}{\Rightarrow x^2 + 2x - 12320 = 0}\) Can you solve this?

Parth (parthkohli):

You'll supposedly get a positive and a negative solution, but you choose the value of x where it is positive. Cheers!

Parth (parthkohli):

Do wolfram for the easy work..

OpenStudy (anonymous):

can you solve it?